Circular No. 7768

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE Further to IAUC 7767, SNe 2001hi, 2001hj, 2001hn, and 2001ho were

present in the Sept. 11 images but were not present on images taken in Oct./Nov. 1999 (cf. IAUC 7312). SN 2001hu was not present on 2001 Oct. 13 (limiting mag 25.4); the remaining supernovae were not present on Sept. 11 (limiting mag 25.8).

SUPERNOVA 2001ib IN NGC 7242 The discovery of an apparent supernova (mag 15.3), located ∼ 9′′–10′′

west and 1′′ north of the center of NGC 7242, has been reported independently on unfiltered CCD images taken by M. Armstrong, Rolvenden, Kent, England, on Dec. 7.808 and 7.844 UT (0.35-m reflector) and by R. Arbour, South Wonston, Hampshire, England, on Dec. 7.786 (0.3-m reflector). Armstrong measures SN 2001ib to be at α = 22^{h}15^{m}38^{s}.72, δ = +37^{o}17′56′′.0 (equinox 2000.0); Arbour provides position end figures 38^{s}.43, 56′′.0. T. Boles obtained confirming images on Dec. 7.824 and 7.828. Nothing is present at this location on Armstrong’s images taken on Aug. 20 (limiting mag 19.5) and Nov. 15 (limiting mag 19.0) or on a red Palomar Sky Survey image taken 1986 Oct. 1 (limiting mag 20.8).

SUPERNOVA 2001hf IN MCG –03-23-17 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2001hf (cf. IAUC 7765), obtained by E. Falco on Dec. 6.51 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova near maximum light. The spectrum consists of a blue continuum with superposed P-Cyg lines of hydrogen. Adopting the NASA/IPAC Extragalactic Database recession velocity of 4486 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 6900 km/s. In addition, images of SN 2001hf obtained by R. Kirshner and S. Jha on Dec. 5.31 with the Baade 6.5-m telescope at Las Campanas yield the following precise position for the supernova: α = 9^{h}05^{m}07^{s}.46, δ = −18^{o}31′04′′.5 (equinox 2000.0; uncertainty ± 0′′.08). Derived magnitudes are R ∼ 16.6 and I ∼ 16.3. The fine seeing conditions (FWHM ∼ 0′′.4) during these observations reveal that a bright foreground star is situated near the nucleus of MCG –03-23-17 and that the nucleus is elongated; the supernova is 4′′.47 west and 6′′.94 north of the star, and 8′′.95 west and 4′′.28 north of the apparent center of the nucleus.

2001 December 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green